OpenStudy (anonymous):

After receiving a discount of 7.5% on its bulk order of notebooks, John's Office Supply pays $7400. What was the price of the order before the discount?

OpenStudy (anonymous):

a. $7955 c. $8000 b. $6845 d. $7215

OpenStudy (anonymous):

@kelliegirl33

OpenStudy (anonymous):

is d right

OpenStudy (anonymous):

oh nvm

OpenStudy (anonymous):

no

OpenStudy (anonymous):

x = original amount before the discount x - 7.5%x = 7400 -- turn percent into decimal x - .075x = 7400 0.925x = 7400 x = 7400/0.925 x = 8000 any questions ?

OpenStudy (anonymous):

no i already got it because i marked out b and d because the were lower